1 General Information: README

This is the README file for the distribution of ESS version 5.7.1

ESS is a GNU Emacs and XEmacs mode for interactive statistical programming and data analysis. Languages supported: the S family (S, S-PLUS and R), SAS, BUGS/JAGS, Stata and XLispStat. ESS grew out of the desire for bug fixes and extensions to S-mode and SAS-mode as well as a consistent union of their features in one package.

Installation instructions are provided in sections for both Unix and Windows; see below. The current development team is led by Martin Maechler since August 2004.

Former project leader A.J. (Tony) Rossini (rossini@blindglobe.net) did the initial port to XEmacs and has been the primary coder. Martin Maechler

(maechler@stat.math.ethz.ch) and Kurt Hornik (Kurt.Hornik@R-project.org)

have assisted with the S family and XLispStat. Stephen Eglen (stephen@gnu.org) has worked mostly on R support. Richard M. Heiberger (rmh@temple.edu) has assisted with S/S-PLUS development for Windows. Richard and Rodney A. Sparapani (rsparapa@mcw.edu) have done much of the work improving SAS batch and interactive support. Rodney has also extended ESS to support BUGS/JAGS and has an interest in improving Stata support.

We are grateful to the previous developers of S-mode (Doug Bates, Ed Kademan, Frank Ritter, David M. Smith), SAS-mode (Tom Cook) and Stata-mode (Thomas Lumley).

1.1 License

ESS is free software; you can redistribute it and/or modify it under the terms of the GNU General Public License as published by the Free Software Foundation; either version 2, or (at your option) any later version.

ESS is distributed in the hope that it will be useful, but WITHOUT ANY WARRANTY; without even the implied warranty of MERCHANTABILITY or FITNESS FOR A PAR-TICULAR PURPOSE. See the GNU General Public License in the file COPYING in the same directory as this file for more details.

1.2 Stability

All released versions (5.x, x >= 4) are meant to be release-quality versions. While some new features are being introduced, we are cleaning up and improving the interface. We know about some remaining documentation inconsistencies. Patches or suggested fixes with bug reports are much appreciated!

1.3 Requirements

ESS is most likely to work with current/recent versions of the following statistical packages: R/S-PLUS, SAS, Stata and JAGS. The one exception is BUGS which is quite old, but should still work.

ESS supports current/recent versions of GNU Emacs and XEmacs. Most notably, the GNU Emacs 22 series and XEmacs 21.4.14 or higher (beta/pre-release versions are NOT SUPPORTED).

1.4 Getting the Latest Version

The latest released version of ESS is always available on the web at: ESS web page

(http://ess.r-project.org) or StatLib (http://lib.stat.cmu.edu/general/ESS/)

The latest development version of ESS is available via https://svn.R-project.org/ESS/, the ESS Subversion repository. If you have a Subversion client (see http://subversion.tigris.org/), you can download the sources using:

      % svn checkout https://svn.r-project.org/ESS/trunk path

which will put the ESS files into directory path. Later, within that directory, ‘svn update’ will bring that directory up to date. Windows-based tools such as TortoiseSVN are also available for downloading the files. Alternatively, you can browse the sources with a web browser at: ESS SVN site (https://svn.r-project.org/ESS/trunk). However, please use a subversion client instead to minimize the load when retrieving.

If you remove other versions of ESS from your emacs load-path, you can then use the development version by adding the following to .emacs:

     (load "/path/to/ess-svn/lisp/ess-site.el")

Note that https is required, and that the SSL certificate for the Subversion server of the R project is

      Certificate information:
       - Hostname: svn.r-project.org
       - Valid: from Jul 16 08:10:01 2004 GMT until Jul 14 08:10:01 2014 GMT
       - Issuer: Department of Mathematics, ETH Zurich, Zurich, Switzerland, CH
       - Fingerprint: c9:5d:eb:f9:f2:56:d1:04:ba:44:61:f8:64:6b:d9:33:3f:93:6e:ad

(currently, there is no “trusted certificate”). You can accept this certificate permanently and will not be asked about it anymore.

1.5 Installation

1.6 Unix installation

For a Unix or Unix-like installation, please follow the next steps. Retrieve the latest tgz file (‘ess-VERSION.tgz’) from ESS downloads area

(http://ess.r-project.org/index.php?Section=download).

GNU Emacs Simple Instructions: for recent versions of Emacs (22.x or higher) and ESS, the installation process is simple.

1. Extract all the files from ‘ess-VERSION.tgz’ when you are in the current working directory of ‘/usr/local/share/emacs/site-lisp’ which exists for packages like ESS (and for Mac OS X create a link to it from the directory

    ‘/Applications/Emacs.app/Contents/Resources/site-lisp’):
         %prompt gtar zxf ess-VERSION.tgz            # for GNU tar
         %prompt gunzip < ess-VERSION.tgz | tar xf - # for Unix tar

2. Then, add the line

         (require 'ess-site)

to ‘~/.emacs’ and restart Emacs.

3. If you see a buffer named ‘*ESS*’, then the simple instructions were most likely successful. If not, then read further.

4. It could be that you have an older version of Emacs, some other problem with your installation or you are not a sysadmin. Whatever the case, you may need to edit ‘ess-site.el’ manually. If that is the case, then you should create a directory just for ESS like ‘~/ess’ or ‘/usr/local/ess’ and unpack ESS there. That way, your changes to ‘ess-site.el’ will not be lost if you update Emacs later.

5. Replace the line above with

     (load "~/ess/ess-VERSION/lisp/ess-site")

in ‘~/.emacs’ and restart Emacs.

6. If you see a buffer named ‘*ESS*’, then the manual instructions were most likely successful. If not, then send a query to ess-help@stat.math.ethz.ch explicitly describing your problem and your environment including operating system, Emacs version, ESS version, etc.

XEmacs Simple Instructions: for recent versions of XEmacs (21.x or higher) and ESS, the installation process is simple.

1. Extract all the files from ‘ess-VERSION.tgz’ when you are in the current

working directory of ‘/usr/local/LOCATION/xemacs/site-packages’ which

exists for packages like ESS where ‘LOCATION’ is ‘lib’ for legacy installations and ‘share’ now (and for Mac OS X create a link to it from the directory

‘/Applications/XEmacs.app/Contents/Resources/site-lisp’):
     %prompt gtar zxf ess-VERSION.tgz            # for GNU tar
     %prompt gunzip < ess-VERSION.tgz | tar xf - # for Unix tar

2. Then, add the line

     (require 'ess-site)

to ‘~/.xemacs/init.el’ and restart XEmacs.

3. If you see a buffer named ‘*ESS*’, then the simple instructions were most likely successful. If not, then read further.

4. It could be that you have an older version of XEmacs, some other problem with your installation or you are not a sysadmin. Whatever the case, you may need to edit ‘ess-site.el’ manually. If that is the case, then you should create a directory just for ESS like ‘~/ess’ or ‘/usr/local/ess’ and unpack ESS there. That way, your changes to ‘ess-site.el’ will not be lost if you update XEmacs later.

5. Replace the line above with

     (load "~/ess/ess-VERSION/lisp/ess-site")

in ‘~/.xemacs/init.el’ and restart XEmacs.

6. If you see a buffer named ‘*ESS*’, then the manual instructions were most likely successful. If not, then send a query to ess-help@stat.math.ethz.ch explicitly describing your problem and your environment including operating system, XEmacs version, ESS version, etc.

1. (OPTIONAL) COMPILING E-LISP:

Edit the default locations of LISPDIR, INFODIR and ETCDIR in Section 1 of ‘Makeconf’ (if you are using XEmacs, then edit the XEmacs subsection in Section 1). You can compile those files by:

     make all

When that completes successfully, install the compiled files:

     make install

1.7 Microsoft Windows installation

For Microsoft Windows installation, please follow the next steps. Retrieve the latest zip file (‘ess-VERSION.zip’) from ESS downloads area

(http://ess.r-project.org/index.php?Section=download).

GNU Emacs Simple Instructions: for recent versions of Emacs (22.x or higher) and ESS, the installation process is simple.

1. Extract all the files from ‘ess-VERSION.zip’ (by double clicking on it and selecting “Extract all files” which launches the Folders Extraction Wizard) into the ‘site-lisp’ directory that exists for packages like ESS. If GNU Emacs was installed in the default location, then this directory can be found somewhere like ‘C:\Program Files\GNU

Emacs\emacs-22.x\site-lisp’.

2. Add the line

     (require 'ess-site)

to ‘~/.emacs’ and restart Emacs.

3. If you see a buffer named ‘*ESS*’, then the simple instructions were most likely successful. If not, then read further.

4. It could be you have an older version of Emacs or some other problem with your instal-

lation. Either way, you may need to edit ‘C:\ess\ess-VERSION\lisp\ess-site.el’

manually. If that is the case, then you should create a directory just for ESS like ‘C:\ess’ and unpack ESS there. That way, your changes to ‘C:\ess\ess-VERSION\lisp\ess-site.el’ will not be lost if you update Emacs later.

5. Replace the line above with

     (load "C:/ess/ess-VERSION/lisp/ess-site")

in ‘~/.emacs’ and restart Emacs.

6. If you see a buffer named ‘*ESS*’, then the manual instructions were most likely successful. If not, then send a query to ess-help@stat.math.ethz.ch explicitly describing your problem and your environment including operating system, Emacs version, ESS version, etc.

XEmacs Simple Instructions: for recent versions of XEmacs (21.x or higher), the installation process is much simpler. Hopefully, these simple instructions will work for you. If not, then more detailed, manual instructions follow.

1. Extract all the files from ‘ess-VERSION.zip’ (by double clicking on it and selecting “Extract all files” which launches the Folders Extraction Wizard) into the ‘site-packages’ directory that exists for packages like ESS. If XEmacs was installed in the default location, then this directory can be found at ‘C:\Program Files\XEmacs\site-packages’.

2. XEmacs requires the HOME environment variable to be defined. You can create it by

visiting the following dialog: My Computer->Control Panel->System->Advanced-

>Environment Variables In the User variables window, press New. And create a variable named HOME with a value something like (you must use forward slashes /

rather than backslashes \) c:/Documents and Settings/%USERNAME%/Application

Data. Then press OK for that window and press OK for the main window. If you also have GNU Emacs installed, GNU Emacs will recognize HOME and expand ~ accordingly.

3. Now launch XEmacs and do a C-x C-f followed by a ~. From the Subdir menu, select Create Directory, and enter .xemacs

4. Add the line

     (require 'ess-site)

to ‘~/.xemacs/init.el’ and restart XEmacs.

5. If you see a buffer named ‘*ESS*’, then the simple instructions were most likely successful. If not, then read further.

6. It could be you have an older version of XEmacs or some other problem with your instal-

lation. Either way, you may need to edit ‘C:\ess\ess-VERSION\lisp\ess-site.el’

manually. If that is the case, then you should create a directory just for ESS like ‘C:\ess’ and unpack ESS there. That way, your changes to

‘C:\ess\ess-VERSION\lisp\ess-site.el’ will not be lost if you update XEmacs

later.

7. Replace the line above with

     (load "C:/ess/ess-VERSION/lisp/ess-site")

in ‘~/.xemacs/init.el’ and restart XEmacs.

8. If you see a buffer named ‘*ESS*’, then the manual instructions were most likely successful. If not, then send a query to ess-help@stat.math.ethz.ch explicitly describing your problem and your environment including operating system, XEmacs version, ESS version, etc.

Now, you should be ready to use ESS. For example, to edit statistical programs, load the files with the requisite extensions (".sas" for SAS, ".S" or "s" or "q" or "Q" for S-PLUS, ".r" or ".R" for R, and ".lsp" for XLispStat). One further step is needed if you wish to run statistical processes, see below.

To run statistical processes under ESS, Windows users will need to make sure that the directories for the software they will be using is in the PATH environment variable.

On Windows NT/2000/XP, add the directories to the PATH using the My Computer-

>Control Panel->System->Advanced->Environment Variables menu. Note that the di-

rectory containing the program is added to the PATH, not the program itself. One such line is needed for each software program. Be sure to use the abbreviation progra~1 and not the long version with embedded blanks as this may cause problems. Also, make sure to use backslashes \ since Windows requires them.

An alternative, for R users, is that rather than adjusting the PATH variable, you can add the following to your emacs initialization file (and restart emacs):

 (setq inferior-R-program-name "c:/progra~1/R/R-2.2.1/bin/Rterm.exe")

This assumes that you have installed R-2.2.1 in the default location. Change the path otherwise to point to other locations.

Windows users who place S-PLUS anywhere other than the default location will also need to add the following three lines (properly adjusted for their location):

     (setq-default inferior-S+6-program-name
         "c:/progra~1/Insightful/SPLUS70/cmd/Splus")
     (setq-default inferior-Sqpe+6-SHOME-name
         "c:/progra~1/Insightful/SPLUS70")
     (setq-default inferior-Sqpe+6-program-name
         "c:/progra~1/Insightful/SPLUS70/cmd/Sqpe.exe")

The above example uses the default location of S-PLUS in c:\progra~1\Insightful. Please note that ESS considers S-PLUS 6, 7, and 8 to be variants of S+6.

These users may also need to modify the emacs variable ess-SHOME-versions to match their installation in order to get the full set of S-PLUS versions on their machine into the ESS menu.

To start the S-PLUS [678].x GUI from ESS under emacs:

1. If you use Cygwin bash as your primary shell, then

         M-x S
         (or M-x S+6).

2. If you use the MSDOS prompt window as your primary shell, then

         M-x S+6-msdos

You will then be asked for a pathname ("S starting data directory?"), from which to start the process. The prompt will propose your current directory as the default. ESS will start the S-PLUS GUI. There will be slight delay during which emacs is temporarily frozen. ESS will arrange for communication with the S-PLUS GUI using the DDE protocol. Send lines or regions from the emacs buffer containing your S program (for example, ‘myfile.s’) to the S-PLUS Commands Window with the C-c C-n or C-c C-r keys. (If you are still using S-PLUS 4.x or 2000,\ then use M-x S+4 or M-x S+4-msdos.)

To start an S-PLUS [678].x session inside an emacs buffer—and without the S-PLUS GUI:

     M-x Sqpe
     (or M-x Sqpe+6).

This works with both the bash and msdos shells. You will then be asked for a pathname ("S starting data directory?"), from which to start the process. The prompt will propose your current directory as the default. You get Unix-like behavior, in particular the entire transcript is available for emacs-style search commands. Send lines or regions from the emacs buffer containing your S program (for example, ‘myfile.s’) to the *S+6* buffer with the C-c C-n or C-c C-r keys. Interactive graphics are available with Sqpe by using the java library supplied with S-PLUS 6.1 and newer releases. Enter the commands:

     library(winjava)
     java.graph()

Graphs can be saved from the java.graph device in several formats, but not PostScript. If you need a PostScript file you will need to open a separate postscript device. (If you are still using S-PLUS 4.x or 2000, then use M-x Sqpe+4.)

To connect to an already running S-PLUS GUI (started, for example, from the S-PLUS icon):

     M-x S+6-existing

or

M-x S+6-msdos-existing

You will then be asked for a pathname ("S starting data directory?"), from which to start the process. The prompt will propose your current directory as the default. ESS will arrange for communication with the already running S-PLUS GUI using the DDE protocol. Send lines or regions from the emacs buffer containing your S program (for example, ‘myfile.s’) to the S-PLUS Commands Window with the C-c C-n or C-c C-r keys. (If you are still using

S-PLUS 4.x or 2000, then use M-x S+4-existing or M-x S+4-msdos-existing.)

If you wish to run R, you can start it with:

M-x R

XLispStat can not currently be run with

M-x XLS

Hopefully, this will change. However, you can still edit with emacs, and cut and paste the results into the XLispStat *Listener* Window under Microsoft Windows.

1.8 Starting an ESS process

To start an S session on Unix or on Windows when you use the Cygwin bash shell, simply

type M-x S RET.

To start an S session on Windows when you use the MSDOS prompt shell, simply type

M-x S+6-msdos RET.

1.9 Current Features

• Languages Supported:

• S family (S 3/4, S-PLUS 3.x/4.x/5.x/6.x/7.x/8.x, and R)

• SAS

• BUGS/JAGS

• Stata

• XLispStat including Arc and ViSta

• Editing source code (S family, SAS, BUGS/JAGS, XLispStat) • Syntactic indentation and highlighting of source code

• Partial evaluation of code

• Loading and error-checking of code

• Source code revision maintenance

• Batch execution (SAS, BUGS/JAGS)

• Use of imenu to provide links to appropriate functions

• Interacting with the process (S family, SAS, XLispStat) • Command-line editing

• Searchable Command history

• Command-line completion of S family object names and file names

• Quick access to object lists and search lists

• Transcript recording

• Interface to the help system

• Transcript manipulation (S family, XLispStat) • Recording and saving transcript files

• Manipulating and editing saved transcripts

• Re-evaluating commands from transcript files

• Help File Editing (R)

• Syntactic indentation and highlighting of source code.

• Sending Examples to running ESS process.

• Previewing

1.10 New Features

Changes/New Features in 5.7:

• ESS[R]: loading a source file (C-c C-l) now works in Windows, similarly to other

platforms; (further; it had accidentally been broken in ESS 5.6 on all platforms) Changes/New Features in 5.6:

• ESS[R]: help() calls have to differ from old default, with newer versions of R;

currently via .help.ESS <- function(...) hack. Changes/New Features in 5.4:

• ESS[SAS]: The long overdue change from make-regexp to regexp-opt for fontlocking is complete. The new regexp-opt is now the default since it is better than the old code in many ways (and especially more maintainable). However, there are certainly some special cases missed (bug reports and patches welcome!). Setting ess-sas-run-regexp-opt to nil will result in the old code being used.

• ESS[BUGS] and ESS[JAGS]: typing = now results in <-.

• ESS[R] function arguments “show” (ess-r-args-show) now uses the new (tooltip-show-at-point) contributed by Erik Iverson.

• Toolbar icons now also work in (beta) Emacs 23.

• ESS[S]: New function ess-change-directory for setting both emacs’ current directory and the directory of an *R* or *S* buffer.

• ESS[S] when transient-mark-mode is true, the mark is now kept, rather than deac-

tivated, thanks to a patch from David Reitter. Changes/New Features in 5.3.11:

• ESS[SAS]: work around bug in Emacs 22.2 & 22.3 which fails to set case-fold fontification automatically.

• Rd mode: support new keyword ’Rdversion’

• ESS[R]: now again works with Emacs 21.x Changes/New Features in 5.3.10:

• Fixed noweb-mode bug accidentally introduced into 5.3.9

• In noweb-mode, e.g., Rnw-mode, electric “<” also inserts closing “@". Further, the code chunk boundaries are better kept up-to-date, such that code[R] <-> text[LaTeX] minor mode switching should happen more reliably.

• In noweb-mode, fix a buglet in rare [Enter] or [Tab] behavior; further, by default disable the former ‘[[’ .. ‘]]’ code-protection-when-filling behavior which has been found to be buggy.

Changes/New Features in 5.3.9:

• ESS[SAS]: evince PDF viewer now supported as well; search order: evince, Xpdf, Adobe/Acrobat Reader

• ESS[R]: added support for Roxygen, potentially to be extended.

• ESS[S] (and R): inferior (*R*) and transcript modes no longer fontify language keywords (such as for, in, etc).

• iESS[Stata]: Customize the ess-sta-delimiter-friendly setting to t to convert embedded semi-colons to newlines for Stata processing.

• Sweave fix for embedded blanks in PDF reader and PDF files

• Several fixes for Major Mode Convention violations in ess-mode and noweb-mode.

• ESS[JAGS]: M-x comment-region now available!

• ESS[S] The ess-swv-* commands (and keybindings) are now in a submenu of the “Noweb” menu, when editing Sweave files.

Changes/New Features in 5.3.8:

• ESS[JAGS]: more separation from ESS[BUGS] (as much as is currently planned); now C-c C-c on an empty ‘.jmd’ creates a template as it should; symbolic links are created for CODA output so BOA is happy: from ‘index.txt’ to ‘.ind’ and

‘chain#.txt’ to ‘#.out’
• ESS[SAS]: buffer-local ess-sas-submit-command and ess-sas-submit-command-
options now recognized by ess-sas-submit-region

• ESS[S]: When trying to evaluate code in an S language buffer and there is no associated process, now start R automatically instead of signalling an error. Also, restart R if there is an associated process which is not running. However, do not start R just via the “electric” ( (ess-r-args-auto-show).

• ESS[S]: For (one-line) functions withOUT ’{ .. }’ bodys, the end of function is now correctly found more often. This notably improves C-c C-c (ess-eval-function-

or-paragraph-and-step).

• ESS[JAGS]: cleanup/re-organization of elisp code; symbolic links for CODA output are now only created by the new JAGS system command in version 1.0.3; specify whether this command is available via ess-jags-system; if not present, then no links are created so that the *shell* buffer does not become unresponsive during the batch run

Changes/New Features in 5.3.7:

• ESS: ess-default-style now *is* customizable, i.e., changing its value in ‘~/.emacs’ now does have the desired effect.

• ESS: ess-font-lock-mode is a new variable (default: t) which controls whether font-locking is enabled in ESS buffers.

• ESS[R]: for XEmacs on Windows; another tweak to find R versions • ESS[SAS]: font-locking updated for ODS and SAS Bayesian Procedures; a more

consistent handling of SAS options by ess-sas-submit-command-options which is buffer-local; portable snooze for MS Windows via customize-able ess-sleep-for (floats welcome); Xpdf now supported as a PDF viewer

• ESS[Rnw]: now also works with “emacs -nw” and Emacs 22.

• ESS[JAGS]: now requires JAGS 1.0 (see the new ESS for JAGS help section for more information): both need work; suggestions welcome

• ESS[R]: [TAB] completion now uses the R-internal completion mechanism (for R >= 2.5.0).

• ESS[R], [S]: interpretation of “ ” as assignment has been removed in

ess-continued-statement-p for R and S.

• several internal code cleanups.

• ESS[R]: An experimental version of a new command Rgui on MS Windows to send lines directly from emacs to Rgui is available in file lisp/essd-rgui.el. Preliminary documentation is in file doc/rgui-doc.txt.

Changes/New Features in 5.3.6:

• ESS: for XEmacs, using “gnuclient” (without a “-q”) works for things like fix() after M-x gnuserv-start has been done.

• ESS[R]: M-x R-newest should now work in more situations on MS Windows, e.g., when R has been installed in a non-default "ProgramFiles" directory tree. In these cases, there’s no need to specify the name (and full path) of the R program anymore.

• ESS[R]: For XEmacs, startup (with new tooltip code) works again.

Changes/New Features in 5.3.5:

• ESS[R] a new defun is available, M-x R-newest, which will start the newest version of R that it can find on your system.

• ESS[R] add Sven Hartenstein’s “R function arguments tips” functionality, via new file ‘../lisp/essd-r-args.el’. Note that this includes an “electric "("” behavior inside R-mode which is active by default and can be customized via ess-r-args-

electric-paren; i.e., use (setq ess-r-args-electric-paren nil) to turn it off.

Further, ess-r-args-show-as allows to switch to the “tooltip” mode.

• ESS: functions ess-get-pdf-viewer and *-ps-viewer; built on new customizable

variables ess-pdf-viewer-pref and ess-ps-viewer-pref; currently used in ess-
swv-PDF and *-PS.

• ESS[R] Improved ess-swv-PDF to run pdf viewer only if pdflatex was ok

• ESS[R] Improved ess-swv-weave to start R automatically if none is running.

• ESS: Do no longer ask which ESS process to use if there is only one.

Changes/New Features in 5.3.4:

• ESS[R] now better work with options(error=recover); and the new default of CHM help files on windows.

• ESS[R] some more cleanup in the “sweave” functions

• miscellaneous fixes Changes/New Features in 5.3.3:

• ESS[S] fix buglet (5.3.2 only) which left command prompt in “execute buffer” and hence help files.

• new customizable variable ess-display-buffer-reuse-frames set to true (which changes default behavior) such that execution or help *frames* are reused.

Changes/New Features in 5.3.2:

• Classic BUGS now supported by (require 'essd-bugs) with ESS[BUGS] and JAGS by (require 'essd-jags) with ESS[JAGS]. But, only one of them can be used at a time since they don’t play nice together. Also, C-c C-c is now bound to the function ess-bugs-next-action (F12 has been retired). And finally, note that ‘essl-bug.el’ is deprecated and the replacement is ‘essl-bugs.el’.

• ESS[R] Improved some of the “Sweave-make” functions (yet scarcely documented)

in ‘ess-swv.el’.

• ESS[S] No longer mess with .Last.value (nor in other “languages”).

Changes/New Features in 5.3.1:

• See the docs for 2 ways to install ESS for XEmacs

1. by uncommenting the XEmacs part of Section 1 of ‘Makeconf’ and performing

    make install

2. by unpacking either ‘ess-5.3.1.tgz’ or ‘ess-5.3.1.zip’ into

    ‘PREFIX/lib/xemacs/site-packages’ on unix or ‘PREFIX\XEmacs\site-packages’

on windows

• ESS[R]: fixed bugs so that Rterm.exe can be found by XEmacs

• ESS[S]: ess-toggle-S-assign-key is slightly changed; in particular, the default

ess-assign-key is now C-x =.

• ESS[R]: M-x R-site-search is a new (slightly experimental) utility similar to R’s RSiteSearch(..) but with the advantage of using Emacs’ preferred browser, see

browse-url-browser-function

Changes/New Features in 5.3.0:

• ESS[BUGS]: sanely re-format statistical output, ‘.bog’, from scientific notation to numbers rounded with 4 decimal places with M-x ess-bugs-sci-round-to-4-dp.

• The keys for navigating among section headings in help buffers worked, but only for one language per session; they should now work for multiple languages. (They were also broken on Windows machines.)

• ESS[S] long standing buglets in the internal logic for loading lisp code on Windows. Particularly fixed behavior in help mode with S-plus GUI.

• New variable, ess-use-inferior-program-name-in-buffer-name, which enables

using the executable name instead of the dialect name for R. Feature request.

• ESS[S] ess-execute-screen-options now also works correctly when there is more than one window *side-by-side* in the same frame and runs in the correct buffer also when there is more than one S buffer.

• iESS[S] new functions ess-eval-paragraph-and-step and ess-eval-function-
or-paragraph-and-step are bound to keys C-c C-p and C-c C-c respectively and

to the menu in ESS-mode; also bound in the help mode (for evaluating examples).

• ESS[S] new function ess-toggle-S-assign-key allows to assign the “ <- ” insertion to an arbitrary key.

1.11 Reporting Bugs

Please send bug reports, suggestions etc. to ESS-bugs@stat.math.ethz.ch The easiest way to do this is within Emacs by typing

 M-x ess-submit-bug-report

This also gives the maintainers valuable information about your installation which may help us to identify or even fix the bug.

If Emacs reports an error, backtraces can help us debug the problem. Type "M-x setvariable RET debug-on-error RET t RET". Then run the command that causes the error and you should see a *Backtrace* buffer containing debug information; send us that buffer.

Note that comments, suggestions, words of praise and large cash donations are also more than welcome.

1.12 Mailing Lists

There is a mailing list for discussions and announcements relating to ESS. Join the list by sending an e-mail with "subscribe ess-help" (or "help") in the body to ess-help-request@stat.math.ethz.ch; contributions to the list may be mailed to ess-help@stat.math.ethz.ch. Rest assured, this is a fairly low-volume mailing list.

The purposes of the mailing list include

• helping users of ESS to get along with it.

• discussing aspects of using ESS on Emacs and XEmacs.

• suggestions for improvements.

• announcements of new releases of ESS.

• posting small patches to ESS.

1.13 Authors

• A.J. Rossini (mailto:blindglobe@gmail.com)
• Richard M. Heiberger (mailto:rmh@temple.edu)
• Kurt Hornik (mailto:Kurt.Hornik@R-project.org)
• Martin Maechler (mailto:maechler@stat.math.ethz.ch)
• Rodney A. Sparapani (mailto:rsparapa@mcw.edu)
• Stephen Eglen (mailto:stephen@gnu.org)